Ask
your own question, for FREE!
Mathematics
10 Online
HELP FAST I WILL FAN AND MEDAL a bag of fruit contains 4 apples, 1 plum, 2 apricots, and 3 oranges. pieces of fruit are drawn twice with replacement what is p( apple, then apricots)?
Still Need Help?
Join the QuestionCove community and study together with friends!
\[A. \frac{ 4 }{ 5 }\] \[B. \frac{ 2 }{25}\] \[C. \frac{ 3 }{ 25 }\] \[D. \frac{ 3 }{ 5}\]
@123.go @animefreakdesu @beccaboo333 @bloopman @ganeshie8 @iPwnBunnies @Jennaaaaaa @mathmale @Preetha @satellite73 @Shay17 @SolomonZelman @Whitemonsterbunny17 @xGuardians @Zale101
@xGuardians
10 Fruits in Total Odds of Apple: 4/10 Odds of Apricot (After Apple): 2/10 4/10 * 2/10 OR (Simplified) 2/5 * 1/5 = 3/25 or 12%
Sorry misread that it said that there IS replacement :)
Still Need Help?
Join the QuestionCove community and study together with friends!
thanxs
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
13 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal